# Continual Learning on CLIP via Incremental Prompt Tuning with Intrinsic Textual Anchors

### v2

@@ 4.4 Ablation Studies
Most importantly, for CIL, during the early incremental stages, models initially encounter only a tiny fraction of the total knowledge to be learned. The model has no idea what it will encounter in the future; thus, learning current data does not take future data into consideration. However, when learning new data, some embedding space is indeed needed to accommodate upcoming tasks (as illustrated in Fig. 1(a)), which causes representation drift (as shown in Fig. LABEL:fig:drift) of previously learned data. The use of textual prototypical contrastive loss helps mitigate this issue by maintaining more stable and meaningful embeddings throughout the incremental learning process.
Effectiveness of Textual Diversity Regularization. Table 4 ablates our diversity loss, which enforces uniform textual prototypes on the unit hypersphere to boost per-class retention. Fig. LABEL:fig:abl_div demonstrates robustness across regularization weights—performance peaks at 0.8, though we adopt 1 by default for simplicity.
Number of prompts. In Fig. LABEL:fig:ps_abl, we evaluate the impact of varying the number of prompts per task on the performance of our methods. We conducted tests with our methods using a range of prompts per task, specifically {1, 2, 3, 4, 5, 10, 20, 50}. Our findings reveal that for TPPT-V on the CUB dataset, using only a single prompt per task is inadequate. Both TPPT-V and TPPT-VT exhibit favorable performance when utilizing at least 5 prompts per task. Consequently, we opt for 10 prompts per task by default, ensuring an effective trade-off between the training budget and performance.
Prompt formulation. In Fig. LABEL:fig:ablate_prompt, we ablate the design of visual and textual prompts utilized in our proposed methods. Overall, both TPPT-V and TPPT-VT show improved performance with deeper prompts, where prompts are embedded into more advanced layers of Transformer blocks. Regarding the length of prompts, extended visual prompts benefit the prediction accuracy for both methods, while the performance of these methods appears relatively insensitive to the length of textual prompts.
Discussion on computation cost. Similar to our design, CODA-P, which also uses incremental visual prompts, introduces 5.68M parameters and reaches an average final accuracy of 70.29. Without the need for additional learnable attention vectors, TPPT-V achieves 78.25% with only 4.3 M parameters, and TPPT-VT further elevates this to 80.07% with an extra 2.46 M parameters.
@@ Acknowledgments
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
@@ Appendix B Additional Experimental Details
We provide additional details of our experiments in this appendix section. Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper.
@@ B.1 Datasets
For a thorough evaluation of our methodologies, we compare the performance with previous methods across a variety of datasets. This includes three popular CIL datasets (Rebuffi et al., 2017; Wang et al., 2022e; d; Smith et al., 2023): CIFAR100 (Krizhevsky et al., 2009), ImageNet-R (Deng et al., 2009), TinyImageNet (Le & Yang, 2015), and CUB200 (Wah et al., 2011), as well as two widely used datasets used for prompt tuning research (Zhou et al., 2022c; Yao et al., 2023; Khattak et al., 2023a; b): FGVCAircraft (Maji et al., 2013) and Stanford Cars (Krause et al., 2013). Each dataset is split into 10 tasks to formulate a CIL setup. We hereby provide a detailed description of these datasets.
@@ B.2 Compared Methods
In our experiments, we conduct a comprehensive comparison of our methods against a variety of previous methods. Firstly, we assess the zero-shot performance of the pre-trained CLIP model (ZS-CLIP) (Radford et al., 2021). This is followed by comparisons with two widely recognized prompt tuning methods: CoOp (Zhou et al., 2022c) and VPT-CL (our implementation of VPT (Jia et al., 2022)), adapted to the CIL scenario. Further, we compare against several popular CL methods: L2P (Wang et al., 2022e), DualPrompt (Wang et al., 2022d), CODA-P (Smith et al., 2023), AttriCLIP (Wang et al., 2023), ZSCL (Zheng et al., 2023), MoE-Adapter (Yu et al., 2024), InfLoRA (Liang & Li, 2024), and PROOF (Zhou et al., 2023c), to provide a comprehensive evaluation of our methods.
@@
VPPT-V (Ours) 13.64 27.61 3.28
VPPT-VT (Ours) 13.11 25.17 3.34
@@
Figure 21: Comparisons to AttriCLIP on CIFAR-100 using ViT-L/14 with pre-trained weight from OpenAI(Radford et al., 2021).
@@
VPPT-V $79.57\pm 0.64$ $80.65\pm 0.29$ $79.98\pm 0.38$ $55.22\pm 0.66$ $95.25\pm 0.81$
VPPT-VT $78.11\pm 0.34$ $80.64\pm 0.67$ $80.33\pm 0.82$ $61.52\pm 0.53$ $94.61\pm 0.59$
Table 9: Average and final accuracy of different methods using OpenAI (Radford et al., 2021) pre-trained weight. All methods are trained with identical exemplar sizes, utilizing the same pre-trained weights, and employing the CLIP model with a ViT-B/16 backbone. $\bar{\mathcal{A}}$ denotes the average of test accuracy over incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and secondary performances are indicated in red and blue, respectively.
@@ Appendix C More Experimental Results and Ablation Studies
Additional comparison with AttriCLIP (Wang et al., 2023). For a fair comparison, we compare our methods with AttriCLIP reported results under the same experiment settings (Wang et al., 2023) on CIFAR-100 in Fig. 21. The backbone model employed was CLIP with a ViT-L/14 architecture, utilizing pre-trained weights from OpenAI. Throughout the incremental learning process, our approach demonstrated superior performance, consistently outperforming AttriCLIP.
@@
Experiment results on Upper-bound performance. In complement to the incremental learning performance in Table 1 and Fig. LABEL:fig:exp_task, we further report the Upper-bound performance of our methods in Table 7, measuring the learning capacity of the model under a non-incremental setting. Notably, for fine-grained datasets such as CUB, Aircraft, and Cars, the performance gap between the final accuracy reported in Table 1 and the Upper-bound performance is marginal, highlighting the effectiveness of the introduction of textual prototypes, which greatly stabilizes the training during the incremental phases.
@@
Studies on Forgetting. In Table 1, we present the average and final accuracies of our proposed methods, alongside comparisons with previous works. Additionally, we evaluate the average forgetting of our methods in Table 5 following (Wang et al., 2022e; d; Smith et al., 2023). We formally denote the forgetting metric for a given task as: $\mathcal{F}_{t}=\frac{1}{t-1}\sum_{\tau=1}^{t-1}\max_{\tau^{\prime}\in\{1,\cdots,t-1\}}(\mathcal{A}_{\tau^{\prime},\tau}-\mathcal{A}_{t,\tau}),$ where $\mathcal{A}_{t,\tau}$ represents the accuracy on task $\tau$ after training on task $t$. We then evaluate average of forgetting over tasks $\tilde{\mathcal{F}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{F}_{t}$. The results demonstrate that our methods not only consistently outperform the previous approaches but also exhibit low levels of forgetting, thus effectively preventing catastrophic forgetting. Benefited from our design of learning with textual prototypes, our method is inherently less prone to forgetting. Enriching textual prototypes with textual diversity not only enhances the stability and efficacy of the training process but also slightly improves the forgetting measure. While CoOp achieves lower forgetting on the Aircraft dataset, it is important to note the substantial difference in prediction accuracy favoring our methods. Given this considerable performance gap, our proposed methods stand out as superior in terms of overall efficacy.
@@
Experiment results with a larger number of tasks. Evaluation results in Table 1 and Fig. LABEL:fig:exp_task were conducted under our default experimental setup, where each dataset is divided into 10 tasks for incremental learning. In Fig. 20, we extend our evaluation to a more challenging experiment setting by splitting each dataset into 20 tasks, which increases the complexity of the learning process and makes the model more prone to forgetting. Despite the increased complexity, our methods, which leverage learning with textual prototypes and the diversity regularization of these prototypes, consistently outperform previous methods in the 20-task split setting.
@@
Computational efficiency of our methods. In addition to the discussion of our parameters in Sec. 4.4, we hereby discuss the computation efficiency of our methods. All experiments are run using one NVIDIA RTX3090 GPU. CODA-P uses a similar design of incremental prompts to us, takes approximately 8ms for training and 5ms for inference per image. TPPT-V takes a comparable or less amount of time, as it does not require additional learnable attention vectors, yet it delivers a significant performance boost of 11% in final accuracy averaged across all datasets compared to CODA-P. TPPT-VT additionally learns textual prompts and further increases final accuracy on the Aircraft dataset by 13%, with only an additional 1.5ms required for both training and inference compared to TPPT-V. While these methods train prompt tokens across multiple layers of the CLIP model, PROOF employs a different approach by training a set of incremental projectors appended to the CLIP encoders, resulting in less processing time. Our methods surpass PROOF by anchoring previous knowledge with textual prototypes and regularizing the learning of multi-modal prompts, thereby enhancing performance without significantly increasing computational overhead.
Experiment results using OpenAI pre-trained CLIP weight. In our main paper, we mainly report the performance using pre-trained weights from LAION-400M (Ilharco et al., 2021). To demonstrate the robustness of our proposed methods, we further evaluate our performance using pre-trained weights from OpenAI (Radford et al., 2021) in Table 9, and we illustrate the test accuracy over incremental stages in Fig. LABEL:supp:openai_exp_fig. Our experiment results demonstrate that our proposed methods are insensitive to the selection of pre-trained weights, consistently outperforming previous methods.

### v3

@@ 4.4 Ablation Studies
Most importantly, for CIL, during the early incremental stages, models initially encounter only a tiny fraction of the total knowledge to be learned. The model has no idea what it will encounter in the future; thus, learning current data does not take future data into consideration. However, when learning new data, some embedding space is indeed needed to accommodate upcoming tasks (as illustrated in Fig. 1(a)), which causes representation drift (as shown in Fig. 4.2) of previously learned data. The use of textual prototypical contrastive loss helps mitigate this issue by maintaining more stable and meaningful embeddings throughout the incremental learning process.
Effectiveness of Textual Diversity Regularization. Table 4 ablates our diversity loss, which enforces uniform textual prototypes on the unit hypersphere to boost per-class retention. Fig. 4.4 demonstrates robustness across regularization weights, performance peaks at 0.8 and we adopt 1 by default for simplicity.
(a) Effect of the textual diversity loss weight on (a) CUB and (b) Aircraft.
(a) Ablation studies on the number of prompts per task for (a) CUB and (b) Aircraft.
Number of prompts. In Fig. 16(a), we evaluate the impact of varying the number of prompts per task on the performance of our methods. We conducted tests with our methods using a range of prompts per task, specifically {1, 2, 3, 4, 5, 10, 20, 50}. Our findings reveal that for TPPT-V on the CUB dataset, using only a single prompt per task is inadequate. Both TPPT-V and TPPT-VT exhibit favorable performance when utilizing at least 5 prompts per task. Consequently, we opt for 10 prompts per task by default, ensuring an effective trade-off between the training budget and performance.
Prompt formulation. In Fig. 16(b), we ablate the design of visual and textual prompts utilized in our proposed methods. Overall, both TPPT-V and TPPT-VT show improved performance with deeper prompts, where prompts are embedded into more advanced layers of Transformer blocks. Regarding the length of prompts, extended visual prompts benefit the prediction accuracy for both methods, while the performance of these methods appears relatively insensitive to the length of textual prompts.
(b) Ablations on multi-modal prompt formulation. Sub-figures (a) and (b) analyze the depth of visual and textual prompts, while (c) and (d) investigate the length of visual and textual prompts.
Discussion on computation cost. Similar to our design, CODA-P, which also uses incremental visual prompts, introduces 5.68M parameters and reaches an average final accuracy of 70.29. Without the need for additional learnable attention vectors, TPPT-V achieves 78.25% with only 4.3 M parameters, and TPPT-VT further elevates this to 80.07% with an extra 2.46 M parameters. More Details in Appendix Table 13.
@@ Acknowledgments
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship from CSIRO and acknowledges the support of the Google PhD Fellowship.
@@ A.1 Algorithm Details
A.2 More Discussions with Related Works
DIKI (diki) and CPP (cpp) tune prompts under primarily intra-visual supervision (calibration in DIKI; prototype steering and routing in CPP), which can leave cross-modal alignment under-constrained. TPPT anchors learning to CLIP text prototypes (fixed/learnable) and optimizes a cross-modal TPCL with text-side diversity, stabilizing the image–text space. CLAP4CLIP (jha2024clap4clip) (adapters with probabilistic logit mixtures) and LoRA/MoE-Adapter variants (inflora; boosting) struggle to maintain vision-text alignment across incremental learning stages, leading to suboptimal knowledge retention. SPU (zhang2024overcoming) selects a subset of important MLP layers via gradient-based scoring and updates them directly, which introduces greater computational overhead than optimizing only a small set of prompt tokens. In contrast, TPPT keeps the backbone frozen and anchors learning to CLIP text prototypes (fixed and learnable), optimizing a cross-modal TPCL with text-side diversity to stabilize the image–text embedding space, and TPPT outperforms these methods (Table 2 and Appendix Table 10).
TreeProbe (zhu2023continual) pairs a fixed CLIP zero-shot head with an exemplar probe and uses tree-style gating to choose between stored exemplars and pre-trained knowledge for open-vocabulary increments. AnytimeCL (zhu2024anytime) emphasizes update scheduling and memory compression to support single-sample, streaming updates. CGIL (frascaroli2024clip) performs generative latent replay by training a VAE in CLIP’s feature space to synthesize past examples. These directions prioritize buffer usage/quality, scheduling/compression, or replay—orthogonal to our method, which targets representation learning via textual anchors and cross-modal alignment.
@@ Appendix B Additional Experimental Details
We provide additional details of our experiments in this appendix section.
@@ B.1 Datasets
For a thorough evaluation of our methodologies, we compare the performance with previous methods across a variety of datasets. This includes three popular CIL datasets (icarl; l2p; dualprompt; codaprompt): CIFAR100 (cifar), ImageNet-R (deng2009imagenet), TinyImageNet (le2015tiny), CUB200 (WahCUB2002011), and DomainNet (peng2019moment) as well as two widely used datasets used for prompt tuning research (coop; kgcoop; maple; promptsrc): FGVCAircraft (aircraft) and Stanford Cars (krause20133d). Each dataset is split into 10 tasks to formulate a CIL setup. We hereby provide a detailed description of these datasets.
@@
DomainNet is a large-scale multi-domain image dataset with roughly 0.6M images spanning 345 categories across six visual domains (clipart, infograph, painting, quickdraw, real, sketch). It is widely used for domain adaptation and continual learning benchmarks to assess scalability across many classes and heterogeneous domains.
@@ B.2 Compared Methods
In our experiments, we conduct a comprehensive comparison of our methods against a variety of previous methods. Firstly, we assess the zero-shot performance of the pre-trained CLIP model (ZS-CLIP) (clip). This is followed by comparisons with two widely recognized prompt tuning methods: CoOp (coop) and VPT-CL (our implementation of VPT (vpt)), adapted to the CIL scenario. Further, we compare against several popular CL methods: L2P (l2p), DualPrompt (dualprompt), CODA-P (codaprompt), AttriCLIP (attriclip), ZSCL (zheng2023preventing), MoE-Adapter (boosting), InfLoRA (inflora), SPU (zhang2024overcoming), DIKI (diki), CPP (cpp), CLAP4CLIP (jha2024clap4clip) and PROOF (proof), to provide a comprehensive evaluation of our methods.
@@
SPU. SPU updates a small, selected subset of backbone weights per task (typically in MLP layers) while freezing the rest. A sparsity mask (e.g., top-k by importance/gradients) identifies the parameters to update, limiting growth and preserving pre-trained knowledge; gradients are computed for full layers, but only the selected weights are updated.
DIKI. DIKI is a parameter-efficient CL method for CLIP that adds lightweight residual modules and uses domain-aware calibration/regularization to reduce interference and retain pre-trained knowledge.
CPP. CPP maintains and steers image class mean prototypes with prompt tuning, and uses query–key routing to select task-specific prompts at inference, operating entirely in the visual space.
CLAP4CLIP. CLAP4CLIP proposes probabilistic fine-tuning of CLIP with adapters plus mixture/logit calibration across tasks; improves stability via uncertainty-aware combination of heads.
@@
TPPT-V (Ours) 13.64 27.61 3.28
TPPT-VT (Ours) 13.11 25.17 3.34
@@
Figure 17: Comparisons to AttriCLIP on CIFAR-100 using ViT-L/14 with pre-trained weight from OpenAI(clip).
@@
TPPT-V $79.57\pm 0.64$ $80.65\pm 0.29$ $79.98\pm 0.38$ $55.22\pm 0.66$ $95.25\pm 0.81$
TPPT-VT $78.11\pm 0.34$ $80.64\pm 0.67$ $80.33\pm 0.82$ $61.52\pm 0.53$ $94.61\pm 0.59$
Table 9: Average and final accuracy of different methods using OpenAI (clip) pre-trained weight. All methods are trained with identical exemplar sizes, utilizing the same pre-trained weights, and employing the CLIP model with a ViT-B/16 backbone. $\bar{\mathcal{A}}$ denotes the average of test accuracy over incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and secondary performances are indicated in red and blue, respectively.
@@
(a) Performance of different methods across incremental stages using pre-trained weights from OpenAI. All methods are trained with the same exemplar size using the same pre-trained weight and backbone model of CLIP with ViT-B/16.
@@ Appendix C More Experimental Results and Ablation Studies
Additional comparison with AttriCLIP (attriclip). For a fair comparison, we compare our methods with AttriCLIP reported results under the same experiment settings (attriclip) on CIFAR-100 in Fig. 17. The backbone model employed was CLIP with a ViT-L/14 architecture, utilizing pre-trained weights from OpenAI. Throughout the incremental learning process, our approach demonstrated superior performance, consistently outperforming AttriCLIP.
Table 10: Comparisons with SPU (zhang2024overcoming).
Methods CIFAR100 CUB Aircraft Stanford Cars TinyImageNet
ZS-CLIP 74.47 65.92 74.77 64.67 33.71 23.01 83.78 76.11 69.55 65.59
SPU 86.22 81.31 80.26 71.98 52.40 42.04 88.12 82.22 79.68 74.89
TPPT-V (Ours) 83.88 77.85 85.32 77.01 62.50 52.75 92.06 87.97 83.18 77.51
TPPT-VT (Ours) 85.72 77.75 84.81 78.33 65.17 61.75 93.28 90.57 82.41 78.08
Additional comparison with SPU (zhang2024overcoming). In Table 10, we reproduce SPU under our experimental setup, using the same memory buffer construction method (iCaRL, consistent with our baseline implementations), task splits, and training budget as in the main experiments. Unlike TPPT, which applies parameter-efficient (PEFT) soft prompts, SPU performs sparse parameter updates directly on the MLP layer, which requires gradient computation over the entire linear layer rather than only a few trainable prompt tokens as in TPPT. Consequently, SPU can potentially allow more adaptation than PEFT modules, as reflected in its higher performance on CIFAR100 (a simpler, low-resolution dataset). On fine-grained datasets with more diverse and complex representations, however, TPPT (especially TPPT-VT) shows significant gains over SPU, suggesting that learning with textual anchors is more effective for continual learning of complex semantics. Moreover, regularized textual prototypes further boost performance to state-of-the-art levels (notably on Aircraft and Cars).
Table 11: Comparisons on DomainNet.
Methods DomainNet
5 Tasks 15 Tasks 30 Tasks 69 Tasks
$\bar{\mathcal{A}}$ $\mathcal{A}_{T}$ $\bar{\mathcal{A}}$ $\mathcal{A}_{T}$ $\bar{\mathcal{A}}$ $\mathcal{A}_{T}$ $\bar{\mathcal{A}}$ $\mathcal{A}_{T}$
ZS-CLIP 63.94 56.66 63.31 56.66 63.66 56.66 63.28 56.66
CoOp 57.45 46.20 55.97 49.24 48.45 47.96 46.38 50.07
VPT-CL 68.70 59.42 70.42 62.78 68.74 61.24 71.84 62.49
CODA-P 68.04 59.09 70.68 61.94 69.72 61.75 70.45 61.25
TPPT-V 72.70 63.73 73.28 63.80 72.81 63.20 74.24 64.69
TPPT-VT 73.70 65.34 73.19 64.08 72.81 62.90 73.78 63.63
Comparisons on larger-scale DomainNet (peng2019moment). We evaluate on DomainNet with 345 classes, splitting the label set into sequences of 5, 15, 30, and 69 tasks. As shown in Table 11, TPPT consistently outperforms prior prompt-based continual learning methods across all task sequence lengths, and the relative advantage is maintained as the number of tasks increases. Average and final accuracies remain stable rather than degrading with longer sequences, indicating that the approach scales to larger class numbers and longer task sequences.
Table 12: Unseen-task zero-shot accuracy across tasks. Deltas are w.r.t. ZS-CLIP at each step (+ improvement, – degradation).
Methods Tasks Seen
1 2 3 4 5 6 7 8 9
ZS-CLIP 66.61 68.14 68.41 69.12 71.44 73.85 75.90 80.15 85.00
CODA-P 28.30 (-38.31) 41.10 (-27.04) 47.33 (-21.08) 47.90 (-21.22) 52.30 (-19.14) 55.80 (-18.05) 62.00 (-13.90) 70.50 (-9.65) 82.60 (-2.40)
TPPT-V 69.49 (+2.88) 72.58 (+4.44) 73.66 (+5.25) 72.57 (+3.45) 73.44 (+2.00) 77.65 (+3.80) 80.47 (+4.57) 82.35 (+2.20) 87.60 (+2.60)
TPPT-VT 45.72 (-20.89) 58.08 (-10.06) 64.44 (-3.97) 69.90 (+0.78) 69.88 (-1.56) 75.45 (+1.60) 78.70 (+2.80) 81.35 (+1.20) 90.40 (+5.40)
Discussions on pre-trained knowledge retention. The retention of pre-trained knowledge during and after continual learning has recently attracted growing attention (zheng2023preventing). Following prior work (zheng2023preventing), we assess general knowledge by measuring zero-shot accuracy on unseen tasks and report these results across the CIFAR-100 continual learning sequence (Table 12). Results show: (a) As continual learning progresses, unseen-task accuracy generally increases because the number of unseen tasks decreases, making classification easier as the number of classes shrinks. (b) For CODA-P, a prompt-tuning CL method related to our incremental prompt framework, zero-shot performance on unseen tasks drops significantly relative to ZS-CLIP (frozen CLIP), indicating severe degradation of general knowledge. (c) For TPPT-V, our static textual-prototype anchors guide CLIP’s representation learning with learnable visual prompts, keeping visual representations clustered around pretrained text prototypes and thus preserving the integrity of the embedding space. We even observe a slight improvement over frozen CLIP, suggesting that visual prompts learned from earlier tasks help the foundation model generalize to unseen tasks. (d) For TPPT-VT, we unfreeze textual prototypes with learnable text prompts to capture more representative linguistic semantics beyond pretrained features. This initially reduces unseen accuracy relative to TPPT-V (especially in the first few tasks) because the text embedding space is being adjusted to learn more representative prototypes. Notably, after Task 6, unseen accuracy slightly surpasses ZS-CLIP, indicating that TPPT-VT benefits from learning generalizable multimodal semantics over a few continual learning stages.
@@ 5 Tasks 15 Tasks 30 Tasks 69 Tasks
Experiment results on Upper-bound performance. In complement to the incremental learning performance in Table 1 and Fig. 8(a), we further report the Upper-bound performance of our methods in Table 7, measuring the learning capacity of the model under a non-incremental setting. Notably, for fine-grained datasets such as CUB, Aircraft, and Cars, the performance gap between the final accuracy reported in Table 1 and the Upper-bound performance is marginal, highlighting the effectiveness of the introduction of textual prototypes, which greatly stabilizes the training during the incremental phases.
@@
Studies on Forgetting. In Table 1, we present the average and final accuracies of our proposed methods, alongside comparisons with previous works. Additionally, we evaluate the average forgetting of our methods in Table 5 following (l2p; dualprompt; codaprompt). We formally denote the forgetting metric for a given task as: $\mathcal{F}_{t}=\frac{1}{t-1}\sum_{\tau=1}^{t-1}\max_{\tau^{\prime}\in\{1,\cdots,t-1\}}(\mathcal{A}_{\tau^{\prime},\tau}-\mathcal{A}_{t,\tau}),$ where $\mathcal{A}_{t,\tau}$ represents the accuracy on task $\tau$ after training on task $t$. We then evaluate average of forgetting over tasks $\tilde{\mathcal{F}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{F}_{t}$. The results demonstrate that our methods not only consistently outperform the previous approaches but also exhibit low levels of forgetting, thus effectively preventing catastrophic forgetting. Benefited from our design of learning with textual prototypes, our method is inherently less prone to forgetting. Enriching textual prototypes with textual diversity not only enhances the stability and efficacy of the training process but also slightly improves the forgetting measure. While CoOp achieves lower forgetting on the Aircraft dataset, it is important to note the substantial difference in prediction accuracy favoring our methods. Given this considerable performance gap, our proposed methods stand out as superior in terms of overall efficacy.
@@
Experiment results with a larger number of tasks. Evaluation results in Table 1 and Fig. 8(a) were conducted under our default experimental setup, where each dataset is divided into 10 tasks for incremental learning. In Fig. 16, we extend our evaluation to a more challenging experiment setting by splitting each dataset into 20 tasks, which increases the complexity of the learning process and makes the model more prone to forgetting. Despite the increased complexity, our methods, which leverage learning with textual prototypes and the diversity regularization of these prototypes, consistently outperform previous methods in the 20-task split setting.
@@
Computational efficiency of our methods. In addition to the discussion of our parameters in Sec. 4.4, we hereby discuss the computation efficiency of our methods. In Table 13, we report per-image training/inference time, peak GPU memory, and estimated FLOPs (GFLOPs/image) under identical hardware, batch size, and data pipelines, comparing TPPT with related continual prompt-tuning methods. All experiments are run using one NVIDIA RTX3090 GPU. CODA-P uses a similar design of incremental prompts to us, takes approximately 8ms for training and 5ms for inference per image. TPPT-V takes a comparable or less amount of time, as it does not require additional learnable attention vectors, yet it delivers a significant performance boost of 11% in final accuracy averaged across all datasets compared to CODA-P. TPPT-VT additionally learns textual prompts and further increases final accuracy on the Aircraft dataset by 13%, with only an additional 1.5ms required for both training and inference compared to TPPT-V.
Table 13: Detailed computation costs. We report per-image train/inference time (ms) and peak GPU memory usage (MiB), and estimated FLOPS (GFLOPs/image).
Methods Train Inference FLOPs
Time Memory Time Memory
ZS-CLIP 2.9 1210.17 33.71
CoOp 3.8 2171.6 3.2 1226.4 38.7
VPT 6.2 5928.0 2.9 1161.6 37.1
L2P 6.9 5963.3 2.9 1472.2 68.3
DualPrompt 7.9 5775.1 2.9 1578.4 68.8
CODA-P 8.2 7974.3 5.4 1653.7 68.1
TPPT-V (Ours) 8.2 7518.1 5.4 1335.6 68.1
TPPT-VT (Ours) 9.7 8821.0 5.9 1344.1 72.7
Experiment results using OpenAI pre-trained CLIP weight. In our main paper, we mainly report the performance using pre-trained weights from LAION-400M (openclip). To demonstrate the robustness of our proposed methods, we further evaluate our performance using pre-trained weights from OpenAI (clip) in Table 9, and we illustrate the test accuracy over incremental stages in Fig. 22(a). Our experiment results demonstrate that our proposed methods are insensitive to the selection of pre-trained weights, consistently outperforming previous methods.
@@
Table 14: Experiment results on SigLIP (zhai2023sigmoid).
Methods CIFAR100 CUB Aircraft
$\bar{\mathcal{A}}$ $\mathcal{A}_{T}$ $\bar{\mathcal{A}}$ $\mathcal{A}_{T}$ $\bar{\mathcal{A}}$ $\mathcal{A}_{T}$
ZeroShot 79.07 70.99 78.71 67.43 51.56 40.74
CoOp 77.97 69.44 75.97 64.29 48.06 39.63
VPT 85.99 77.90 80.23 69.80 51.51 46.28
CODA-P 86.69 79.42 86.11 80.36 71.33 69.43
TPPT-V (Ours) 89.16 81.71 88.18 82.19 72.00 69.91
TPPT-VT (Ours) 88.99 81.64 87.90 82.40 74.81 71.08
Generalization to other VLMs. To further validate effectiveness, we evaluate TPPT on SigLIP (zhai2023sigmoid), whose pretraining differs from CLIP and yields stronger zero-shot performance. Despite these differences, TPPT, guided by stable textual anchors, consistently improves the SigLIP backbone across diverse downstream tasks (coarse-grained CIFAR100 and fine-grained CUB and Aircraft), achieving state-of-the-art results relative to other continual prompt tuning methods.